OpenStudy (anonymous):

Derivative of 3lnx/-2x^-1

OpenStudy (raffle_snaffle):

Instead of using quotient rule I suggest you product instead

OpenStudy (anonymous):

\[\frac{ 3\ln \left| x \right| }{ -2x ^{-1}}\]

OpenStudy (raffle_snaffle):